OpenStudy (anonymous):

Use tiles to factor 2x^2 + 10x + 8

OpenStudy (anonymous):

what's tiles factor?

OpenStudy (dumbcow):

=2(x^2 +5x +4) 4*1=4 4+1=5 =2(x+4)(x+1)